What will be the pH of a solution containing 100 mol of potassium hydroxide and 0.05 mol of sulfuric acid in 100 cm3 of solution?


1
Expert's answer
2021-05-21T02:33:40-0400

The reaction can be represented as follows:

2KOH + H2SO4 = K2SO4 + 2H2O

From here, after the reaction, 1.00 mol - (2 × 0.05 mol) = 0.9 mol of KOH are left in the solution.

The molar concentration of KOH equals:

M(KOH) = n(KOH) / V(KOH) = 0.9 mol / 100 cm3 = 0.9 mol / 0.1 L = 9 M

From here, the pOH of the solution:

pOH = −log[OH] = −log[9] = 0.95

Finally, pH equals:

pH = 14 - pOH = 14 - 0.95 = 13.05


Answer: 13.05